# Semiclassics for the QCD vacuum structure through $T^{2}$-compactification with the baryon-’t Hooft flux

### v1

@@ 1 Introduction
Recently, a novel semiclassical approach to exploring confining vacua has been proposed Tanizaki:2022ngt . This semiclassical framework is constructed on $\mathbb{R}^{2}\times T^{2}$ at a small torus $T^{2}$ with ’t Hooft flux (combined with flavor or $U(1)_{B}$ twist if necessary). This compactification with ’t Hooft flux preserves the ’t Hooft anomalies, so we boldly conjecture that the compactified theory is adiabatically connected to the gauge theories on $\mathbb{R}^{4}$ (see Fig. 1). This enables the semiclassical calculation of the confining vacua at least on qualitative aspects. This method has been applied to a variety of confining gauge theories, and it successfully provides reasonable pictures of these theories Tanizaki:2022ngt ; Tanizaki:2022plm ; Hayashi:2023wwi , including $SU(N)$ Yang-Mills theory, $\mathcal{N}=1$ super-Yang-Mills theory, QCD with fundamental quarks, QCD with 2-index quarks, and bifundamental QCD. While the semiclassical approach relies on the nontrivial working assumption of the adiabatic continuity, the resulting consequences align well with our current understanding of confining vacua.
@@
As an application, we revisit the phase diagram of $N_{f}=1+1$ QCD (and $N_{f}=1+1+1$ QCD with a fixed strange mass) on the $(m_{u},m_{d})$ plane, examining it through our semiclassical method. The physics of $N_{f}=1+1$ and $N_{f}=1+1+1$ QCD on the $(m_{u},m_{d})$ plane has gathered interest to deepen our understanding of chiral symmetry breaking and also in the context of the strong $CP$ problem. An interesting phenomenon in this study is the appearance of a spontaneously $CP$-broken phase, known as the Dashen phase Dashen:1970et . This phase is expected to appear in certain regions where $m_{u}m_{d}<0$ with $\theta=0$, which is equivalent to $\theta=\pi$ with both $m_{u},m_{d}>0$.
Previous investigations into this topic basically use the phenomenological chiral Lagrangian, both with and without an (ad-hoc) $\eta^{\prime}$ contribution Creutz:1995wf ; Creutz:2003xc ; Creutz:2003xu ; Creutz:2013xfa ; Aoki:2014moa . Building upon the research, this paper aims to provide a fresh perspective on the phase diagram based on our semiclassical framework. Compared to the conventional chiral Lagrangian, our framework explicitly derives the global structure of $\eta^{\prime}$ degrees of freedom from the microscopic calculation. We note that previous studies give confusing results on the global structure of the phase diagram depending on the choice of the $\eta^{\prime}$ mass term. We will see how our improvement of the $\eta^{\prime}$ field settles the subtle issues on the phase structure in previous studies.
@@ 1.1 Summary and implications
    $~{}~{}~{}~{}~{}~{}~{}~{}~{}-2K\mathrm{e}^{-S_{I}/N}\cos\left(\frac{\eta^{\prime}-\theta}{N}\right),.$ (35)
    where $m$ is the quark mass, $\mu$ is a scale introduced by the bosonization scheme, and $K\mathrm{e}^{-S_{I}/N}$ is the weight of the fractional instanton. When $N_{f}=1$, we just need to forget $\varphi_{1},\ldots,\varphi_{N_{f}-1}$. If we adopt the non-Abelian bosonization, we obtain the 2d analog of the chiral Lagrangian (4.1).
@@
  • When $\eta^{\prime}$ is included, its chiral Lagrangian conventionally has an ambiguity about the $\eta^{\prime}$ mass term; two common options are the Kobayashi-Maskawa-’t Hooft determinant vertex, $(\mathrm{e}^{-\mathrm{i}\theta}\operatorname{det}U+\mathrm{c.c.})$ Kobayashi:1970ji ; Kobayashi:1971qz ; Maskawa:1974vs ; tHooft:1976rip , and the log-det vertex, $(\mathrm{i}\ln\det(U)+\theta)^{2}$, supported by the large-$N$ analysis Witten:1980sp ; DiVecchia:1980yfw ; Sakai:2004cn . The global structure of the phase diagram depends on the choice, and results of previous studies contain some confusions. As a related problem, if we naively add the $\eta^{\prime}$ meson to the chiral Lagrangian, such a model has struggled to reconcile with the discrete anomalies of Refs. Gaiotto:2017tne ; Tanizaki:2017mtm ; Tanizaki:2018wtg ; Anber:2019nze , particularly with the quenched limit.
@@
  • In our approach, these issues are resolved by endowing the $\eta^{\prime}$ field with an extended periodicity of $2\pi N$, $\eta^{\prime}\sim\eta^{\prime}+2\pi N$, and this is a consequence of integrating out the gluon’s $\mathbb{Z}_{N}$ center contribution. We can naturally fix the form of the $\eta^{\prime}$ mass by the symmetry, locality, and dimensional counting, which proposes the improvement of $4$d chiral Lagrangian (40). As an application, we revisit and refine the existence of the Dashen phase, deepening our understandings of these global aspects.
2 Adiabatic continuity, semiclassics, and center vortices
@@
This paper aims to explore the QCD vacuum structure and related physics using the semiclassical framework based on ’t Hooft flux $T^{2}$ compactification, as proposed in Ref. Tanizaki:2022ngt . We here concisely summarize the underlying methodology.
@@
This semiclassical method is predicated on the adiabatic continuity conjecture, which is motivated by the volume independence conjecture. We would like to investigate the physics of confining vacua in $\mathbb{R}^{4}$ from a weakly-coupled theory on $\mathbb{R}^{2}\times T^{2}$ with a small torus $T^{2}$. At small $T^{2}$, the theory becomes weakly-coupled and is described by a 2d low-energy effective theory. Here, quark confinement is explained by the dilute gas approximation of fractional instantons, namely center vortices. The center-vortex picture is one of the popular scenarios in the understanding of quark confinement Greensite:2011zz , which is another advantage of this semiclassical framework in addition to its calculability. (For a similar approach with the dilute gas approximation of monopoles on $\mathbb{R}^{3}\times S^{1}$, see, e.g., Refs. Unsal:2007vu ; Unsal:2007jx ; Unsal:2008ch ; Shifman:2008ja ; Poppitz:2012sw ; Poppitz:2012nz .)
2.1 Main ansatz: adiabatic continuity
The main ansatz for our study is the adiabatic continuity conjecture for the ’t Hooft flux $T^{2}$ compactification. Figure 1 provides a summary of the methodology. This adiabatic continuity for the $T^{2}$ compactification with nontrivial ’t Hooft flux tHooft:1979rtg ; tHooft:1981sps is a highly nontrivial conjecture, but it has been underpinned by theoretical and empirical observations.
One of the observations is that the $T^{2}$ compactification with the ’t Hooft flux preserves the anomaly related to the center symmetry Tanizaki:2017qhf ; Yamazaki:2017dra ; Tanizaki:2022ngt . The anomaly-preservation is a necessary condition for realizing the confinement phase, i.e., preserving the center symmetry, at small $T^{2}$. Note that the preservation of the center symmetry is a criterion for the validity of the large-$N$ volume independence Eguchi:1982nm , and introducing the ’t Hooft twist is one of common techniques in this context GonzalezArroyo:1982hz ; GonzalezArroyo:2010ss .
Empirically, this semiclassical method successfully gives a reasonable picture for confining vacua in several 4d non-Abelian gauge theories: $SU(N)$ Yang-Mills theory, $SU(N)$ $\mathcal{N}=1$ super-Yang-Mills theory, QCD with fundamental quarks, QCD with 2-index quarks, and bifundamental QCD Yamazaki:2017ulc ; Tanizaki:2022ngt ; Tanizaki:2022plm ; Hayashi:2023wwi . This paper expands on the semiclassical analysis for QCD with fundamental quarks proposed in Tanizaki:2022ngt .
2.2 Semiclassics for $SU(N)$ Yang-Mills theory
@@
Here, let us review the semiclassical analysis of the $SU(N)$ Yang-Mills theory, and see how the dilute gas approximation of center vortices gives the confining vacuum. We consider the $SU(N)$ Yang-Mills theory on $\mathbb{R}^{2}\times T^{2}$ at a small torus $T^{2}$, and we take the ’t Hooft boundary condition for $T^{2}$.
@@
Throughout this paper, we use the following coordinates: $x=(x_{1},x_{2},x_{3},x_{4})=(\vec{x},x_{3},x_{4})\in\mathbb{R}^{2}\times T^{2}$, and we identify $(x_{3},x_{4})\sim(x_{3}+L,x_{4})\sim(x_{3},x_{4}+L)$ to form $T^{2}$. The unit ’t Hooft flux can be inserted by introducing the transition functions $g_{3}(x_{4})$ and $g_{4}(x_{3})$ for $T^{2}$ satisfying
$g_{3}(L)^{\dagger}g_{4}(0)^{\dagger}g_{3}(0)g_{4}(L)=\mathrm{e}^{\frac{2\pi\mathrm{i}}{N}}.$ (1)
@@
with $C\propto\operatorname{diag}(1,\mathrm{e}^{\frac{2\pi\mathrm{i}}{N}},\cdots,\mathrm{e}^{\frac{2\pi\mathrm{i}(N-1)}{N}})$ and $(S)_{ij}\propto\delta_{i+1,j}$.
We shall show that this setup provides the confining vacuum even at small $T^{2}$. The important observations are listed as follows Yamazaki:2017ulc ; Tanizaki:2022ngt .
@@
    where $\mathbb{Z}_{N}^{[0]}\times\mathbb{Z}_{N}^{[0]}$ is the conventional center symmetry acting on the Polyakov loops of $T^{2}$. For the adiabatic continuity, the center symmetry $\mathbb{Z}_{N}^{[0]}\times\mathbb{Z}_{N}^{[0]}$ should not be spontaneously broken: the Polyakov loops along $x_{3}$ and $x_{4}$ directions should vanish $\braket{\operatorname{tr}P_{3}}=\braket{\operatorname{tr}P_{4}}=0$. With a periodic boundary condition, the center symmetry is spontaneously broken at a small compactification scale and the adiabatic continuity cannot be achieved. The ’t Hooft boundary condition avoids this problem. Indeed, we can suppose that local fluctuations are negligible on a sufficiently small torus, and the transition functions give the Polyakov loops. In the gauge (2), their classical values are
@@
    With the choice of the transition functions (2), the boundary conditions for gluons are given by
@@
    which admits no constant mode. Thus, the gluon field acquires a Kalzua-Klein mass of order $O(1/L)$ (or more precisely $O(1/NL)$) without zero modes.
@@
    $Q_{\mathrm{top}}=\frac{1}{8\pi^{2}}\int\operatorname{tr}\left((\tilde{f}-B)\wedge(\tilde{f}-B)\right)\in-\frac{1}{N}+\mathbb{Z},$ (6)
@@
    where $S_{I}=8\pi^{2}/g^{2}$ is the instanton action, and the equality holds if and only if the $f$ is (anti-)self-dual. Some numerical studies Gonzalez-Arroyo:1998hjb ; Montero:1999by ; Montero:2000pb support that the anti-self-dual solution indeed exists and this fractional instanton will survive as a local solution in the decompactification limit of $(x_{1},x_{2})$ directions. Thus, we assume that such a solution saturating the BPS bound exists locally, which we call center vortex or fractional instanton. Similar fractional instantons in ×$\mathbb{R}\times T^{3}$ are also numerically studied in GarciaPerez:1989gt ; GarciaPerez:1992fj ; Itou:2018wkm , which are also used in the different but deeply-related semiclassics Yamazaki:2017ulc . For a recent review of the fractional instanton picture of confinement, see also Gonzalez-Arroyo:2023kqv .
Based on these observations, the dilute gas approximation of center vortices provides a reasonable picture of the confining vacuum. The center-vortex and anti-center-vortex amplitudes can be expressed as

### v3

@@ 1 Introduction
Recently, a novel semiclassical approach to exploring confining vacua has been proposed Tanizaki:2022ngt . This semiclassical framework is constructed on $\mathbb{R}^{2}\times T^{2}$ at a small torus $T^{2}$ with ’t Hooft flux tHooft:1979rtg ; tHooft:1981sps (combined with flavor or $U(1)_{B}$ twist if necessary), which maintains ’t Hooft anomalies of $4$d theories in the $2$d effective theory Tanizaki:2017qhf ; Yamazaki:2017dra ; Tanizaki:2022ngt . This $2$d effective theory gives the weak-coupling description of the confinement-related phenomena, and it successfully offers a reasonable picture for various confining theories Tanizaki:2022ngt ; Tanizaki:2022plm ; Hayashi:2023wwi , including $SU(N)$ Yang-Mills theory, $\mathcal{N}=1$ super-Yang-Mills theory, QCD with fundamental quarks, QCD with 2-index quarks, and bifundamental QCD. Based on these observations, let us assume the adiabatic continuity conjecture in this paper: the $2$d semiclassical theory captures qualitative features of $4$d strongly-coupled theories. Figure 1 provides a summary of this methodology.
@@
As an application, we revisit the phase diagram of $N_{f}=1+1$ QCD (and $N_{f}=1+1+1$ QCD with a fixed strange mass)111 In this paper, we use the jargon $N_{f}=1+1$ or $(1+1)$-flavor when the up and down quarks have different masses. Similarly, $N_{f}=1+1+1$ QCD indicates that up, down, and strange masses are all different. on the $(m_{u},m_{d})$ plane, examining it through our semiclassical method. The physics of $N_{f}=1+1$ and $N_{f}=1+1+1$ QCD on the $(m_{u},m_{d})$ plane has gathered interest to deepen our understanding of chiral symmetry breaking and also in the context of the strong $CP$ problem. An interesting phenomenon in this study is the appearance of a spontaneously $CP$-broken phase, known as the Dashen phase Dashen:1970et . This phase is expected to appear in certain regions where $m_{u}m_{d}<0$ with $\theta=0$, which is equivalent to $\theta=\pi$ with both $m_{u},m_{d}>0$.
Previous investigations into this topic basically use the phenomenological chiral Lagrangian, both with and without an $\eta^{\prime}$ contribution Creutz:1995wf ; Creutz:2003xc ; Creutz:2003xu ; Creutz:2013xfa ; Aoki:2014moa . When $\eta^{\prime}$ is included, however, the chiral Lagrangian has an ambiguity about its mass term; two conventional options are the Kobayashi-Maskawa-’t Hooft determinant vertex, $(\mathrm{e}^{-\mathrm{i}\theta}\operatorname{det}U+\mathrm{c.c.})$ Kobayashi:1970ji ; Kobayashi:1971qz ; Maskawa:1974vs ; tHooft:1976rip , and the log-det vertex, $(\mathrm{i}\ln\det(U)+\theta)^{2}$, supported by the large-$N$ analysis Witten:1980sp ; DiVecchia:1980yfw ; Sakai:2004cn . The global aspects of the phase diagram depend on the choice of the $\eta^{\prime}$ mass term, which has led to some confusion. As a related problem, if we naively add the $\eta^{\prime}$ meson to the chiral Lagrangian, such a model has struggled to reconcile with existing knowledge about the global structure, including the discrete anomalies of Refs. Gaiotto:2017tne ; Tanizaki:2017mtm ; Tanizaki:2018wtg ; Anber:2019nze and the vacuum structure at the quenched limit ($m_{u},m_{d}\to\infty$).
This paper aims to provide a fresh perspective on these topics based on our semiclassical framework. Compared to the conventional chiral Lagrangian, our framework explicitly derives the global structure of $\eta^{\prime}$ degrees of freedom from the microscopic calculation. We will see how our improvement of the $\eta^{\prime}$ field settles the subtle issues on the phase structure in previous studies.
@@ 1.1 Summary and implications
    $~{}~{}~{}~{}~{}~{}~{}~{}~{}-2K\mathrm{e}^{-S_{I}/N}\cos\left(\frac{\eta^{\prime}-\theta}{N}\right),$ (38)
    where $m$ is the quark mass, $\mu$ is a scale introduced by the bosonization scheme, and $K\mathrm{e}^{-S_{I}/N}$ is the weight of the fractional instanton. If we adopt the non-Abelian bosonization, we obtain the 2d analog of the chiral Lagrangian (4.1).
@@
  • As mentioned above, the chiral Lagrangian suffers from an ambiguity when we try to introduce the $\eta^{\prime}$ field. In our approach, this issue is resolved by clarifying that the $\eta^{\prime}$ field has an extended periodicity of $2\pi N$, $\eta^{\prime}\sim\eta^{\prime}+2\pi N$, which is a consequence of integrating out the gluon’s $\mathbb{Z}_{N}$ center contribution. We can naturally fix the form of the $\eta^{\prime}$ mass by the symmetry, locality, and dimensional counting, which proposes the improvement of $4$d chiral Lagrangian (43). As an application, we revisit and refine the existence of the Dashen phase, deepening our understanding of these global aspects.
2 Semiclassics for $SU(N)$ Yang-Mills theory: dilute gas of center vortices
@@
This paper aims to explore the QCD vacuum structure and related physics using the semiclassical framework based on ’t Hooft flux $T^{2}$ compactification, as proposed in Ref. Tanizaki:2022ngt . We here concisely summarize the methodology through the example of $SU(N)$ Yang-Mills theory.
@@
Based on the adiabatic continuity conjecture, we investigate the physics of confining vacua in $\mathbb{R}^{4}$ from a weakly-coupled theory on $\mathbb{R}^{2}\times T^{2}$ with a small torus $T^{2}$ (Figure 1). At small $T^{2}$, the theory becomes weakly-coupled and is described by a 2d low-energy effective theory. Here, quark confinement is explained by the dilute gas approximation of fractional instantons, namely center vortices. The center-vortex picture is one of the plausible scenarios to explain quark confinement Greensite:2011zz , and this semiclassical framework gives its explicit and calculable realization through the $T^{2}$ compactification. (For a similar approach with the dilute gas approximation of monopoles on $\mathbb{R}^{3}\times S^{1}$, see, e.g., Refs. Unsal:2007vu ; Unsal:2007jx ; Unsal:2008ch ; Shifman:2008ja ; Poppitz:2012sw ; Poppitz:2012nz .)
@@
Let us review the semiclassical analysis of the $SU(N)$ Yang-Mills theory, and see how the dilute gas approximation of center vortices yields the confining vacuum Tanizaki:2022ngt . We consider the $SU(N)$ Yang-Mills theory on $\mathbb{R}^{2}\times T^{2}$ at a small torus $T^{2}$, and we take the ’t Hooft boundary condition for $T^{2}$.
@@
Throughout this paper, we use the following coordinates: $x=(x_{1},x_{2},x_{3},x_{4})=(\vec{x},x_{3},x_{4})\in\mathbb{R}^{2}\times T^{2}$, and we identify $(x_{3},x_{4})\sim(x_{3}+L,x_{4})\sim(x_{3},x_{4}+L)$ to form $T^{2}$. The unit ’t Hooft flux can be inserted by introducing the $SU(N)$-valued transition functions $g_{3}(x_{4})$ and $g_{4}(x_{3})$ for $T^{2}$ satisfying
$g_{3}(L)^{\dagger}g_{4}(0)^{\dagger}g_{3}(0)g_{4}(L)=\mathrm{e}^{\frac{2\pi\mathrm{i}}{N}}{\color[rgb]{0,0,0}1_{N\times N}}.$ (1)
@@
with $C=\mathrm{e}^{\mathrm{i}\alpha}\operatorname{diag}(1,\mathrm{e}^{\frac{2\pi\mathrm{i}}{N}},\cdots,\mathrm{e}^{\frac{2\pi\mathrm{i}(N-1)}{N}})$, $(S)_{ij}=\mathrm{e}^{\mathrm{i}\alpha}\delta_{i+1,j}$, and $\mathrm{e}^{\mathrm{i}N\alpha}=(-1)^{N+1}$.
We shall show that this setup provides the confining vacuum even at small $T^{2}$. The important observations are listed as follows Yamazaki:2017ulc ; Tanizaki:2022ngt ; Cox:2019aji .
@@
    where $\mathbb{Z}_{N}^{[0]}\times\mathbb{Z}_{N}^{[0]}$ is the conventional center symmetry acting on the Polyakov loops of $T^{2}$. From the adiabatic continuity, the center symmetry $\mathbb{Z}_{N}^{[0]}\times\mathbb{Z}_{N}^{[0]}$ should not be spontaneously broken: the Polyakov loops along $x_{3}$ and $x_{4}$ directions should vanish $\braket{\operatorname{tr}P_{3}}=\braket{\operatorname{tr}P_{4}}=0$. With a periodic boundary condition, the center symmetry is spontaneously broken at a small compactification scale and the adiabatic continuity cannot be achieved222 Without the ’t Hooft flux, the classical vacua are parametrized by the Polyakov loops $P_{3},P_{4}$ with $P_{3}P_{4}=P_{4}P_{3}$, and the one-loop calculation produces the effective potential Gross:1980br ; Luscher:1982ma ; Unsal:2010qh ; Tanizaki:2022ngt , $V[P_{3},P_{4}]=-\frac{2}{\pi^{2}L^{4}}\sum_{(n_{3},n_{4})\neq(0,0)}\frac{1}{(n_{3}^{2}+n_{4}^{2})^{2}}\operatorname{tr}\left(P_{3}^{n_{3}}P_{4}^{n_{4}}\right).$ This has $N^{2}$ minima located at $\frac{1}{N}\braket{\operatorname{tr}P_{3}}=\mathrm{e}^{\frac{2\pi m_{3}}{N}},~{}~{}\frac{1}{N}\braket{\operatorname{tr}P_{4}}=\mathrm{e}^{\frac{2\pi m_{4}}{N}}$, where $m_{3},m_{4}=0,\cdots,N-1$, and $\mathbb{Z}_{N}^{[0]}\times\mathbb{Z}_{N}^{[0]}$ is spontaneously broken. . The ’t Hooft boundary condition avoids this problem. Indeed, we can suppose that local fluctuations are negligible on a sufficiently small torus, and the transition functions give the Polyakov loops. In the gauge (2), their classical values are
@@
    With the choice of the transition functions (2), the boundary conditions for the gauge field are given by
@@
    which admits no constant mode. Thus, the gluon field acquires a Kalzua-Klein mass of order $O(1/NL)$ without zero modes.
@@
    $Q_{\mathrm{top}}=\frac{1}{8\pi^{2}}\int_{{\color[rgb]{0,0,0}T^{2}\times T^{2}}}\operatorname{tr}\left((\tilde{f}-B)\wedge(\tilde{f}-B)\right)\in-\frac{1}{N}+\mathbb{Z},$ (6)
@@
    where $S_{I}=8\pi^{2}/g^{2}$ is the instanton action, and the equality holds if and only if the $f$ is (anti-)self-dual. Some numerical studies Gonzalez-Arroyo:1998hjb ; Montero:1999by ; Montero:2000pb support that the anti-self-dual solution indeed exists and this fractional instanton will survive as a local solution in the decompactification limit of $(x_{1},x_{2})$ directions. Based on these numerical observations, we assume that such a solution saturating the BPS bound with $Q_{\mathrm{top}}=\pm 1/N$ exists locally in our $\mathbb{R}^{2}\times T^{2}$ setup333 Here, “exists locally” means that the solution exists if the boundary conditions at infinity in $\mathbb{R}^{2}$ are not considered. Note that such a solution with fractional topological charge does not exist globally when the point at infinity of $\mathbb{R}^{2}$ is regular, i.e., when $\mathbb{R}^{2}$ can be compactified without any twist. Thus, in order to compute the partition function by the dilute gas approximation, we will consider ensembles of center vortices and anti-vortices, provided that the total topological charge is an integer. , which we call center vortex or fractional instanton444 About the terminology, the “fractional instanton” would be obvious as it has the $\frac{1}{N}$ fractional topological charge compared with the minimal instanton on $S^{4}$. On the other hand, the “center vortex” may not be obvious as it should be defined by the following feature: When the vortex goes across the Wilson loop, its phase is rotated by the $\mathbb{Z}_{N}$ center element. In the $T^{2}$-compactified setup with the ’t Hooft flux, it turns out that the fractional instanton describes the center vortex Gonzalez-Arroyo:1998hjb ; Montero:1999by ; Montero:2000pb , so we identify them in this paper. . Similar fractional instantons in ×$\mathbb{R}\times T^{3}$ are also numerically studied in GarciaPerez:1989gt ; GarciaPerez:1992fj ; Itou:2018wkm , which are also used in the different but deeply-related semiclassics Yamazaki:2017ulc ; Cox:2019aji . For a recent review of the fractional instanton picture of confinement, see also Gonzalez-Arroyo:2023kqv .
Based on these observations, the dilute gas approximation of center vortices provides a reasonable picture of the confining vacuum. We consider ensembles of center vortices with $S=S_{I}/N,~{}Q_{\mathrm{top}}=1/N$ and anti-vortices with $S=S_{I}/N,~{}Q_{\mathrm{top}}=-1/N$. The center-vortex and anti-center-vortex amplitudes can be expressed as